# Quantum Theory of Optical Spin Texture in Chiral Tellurium Lattice

### v2

@@ Abstract
The absence of inversion symmetry in chiral tellurium (Te) creates unique spin textures within its electron waves. However, understanding textured optical waves within Te remains a challenge due to the semi-classical limitations of long-wavelength approximation. To unveil textured optical eigenwaves in Te, we develop a spin-resolved deep-microscopic optical bandstructure analogous to its electronic counterpart. We demonstrate that the degeneracies in this optical bandstructure is lifted by the twisted lattice of Te, which induces optical gyrotropy. Our theory shows excellent agreement with experimental optical gyrotropy measurements. At the lattice level, we reveal that the chirality of Te manifests as deep-microscopic optical spin texture within the optical wave. Our framework uncovers the finite-momentum origin of optical activity and provides a microscopic basis for light-matter interactions in chiral crystalline materials.
@@
Figure 2: Super-dispersive optical gyrotropy and chirality density for undoped Te. (a) Experimental plot of the projection of the longitudinal gyration tensor component $\bm{g}_{33}$ on the photon momentum, vs. (b) Theoretical prediction using Purdue-PicoMax at $\bm{q}=(0,0,0.1)2\pi/a$ The Feynman diagram in the inset reflects the transverse SILC of the induced transverse current including the momentum-exchange processes. Here, $c$ and $v$ represent the conduction and valence bands respectively. (c,d) The spin-resolved chirality density projection and total spin on the isofrequency surface of RCP ($+$) at $\omega=60$THz (c) and LCP (--) (d). The concentration near the pole of the sphere and the difference between RCP and LCP illustrates the suitable axis (optic axis) for observing high optical activity.
@@
Figure 4: Deep-microscopic optical spin texture reveal orientation of polarization of optical waves. Vectorial distribution of the optical spin texture of the optical waves in the first transverse optical band. The colormap shows the projection of the spin density along the screw axis. The plots are shown for three different planes of the Te crystal: (a) $x=0$, (b) $z=0$, and (c) $y=0$.
Our work reveals that the lack of inversion symmetry in Te lifts the degeneracy in the optical bandstructure, giving rise to optical gyrotropy. The deep-microscopic optical bandstructure captures the super-dispersive nature of this optical gyrotropy, and shows excellent agreement with experimental observations. We have implemented the necessary tools for calculating this optical gyrotropy within our open-source software: Purdue-PicoMax [8]. Previously, there have been attempts to compute the optical gyrotropy through first-principles calculations [9, 10, 11, 12, 13]. However, they do not quantitatively capture the experimentally measured dispersion of optical gyrotropy across all wavelengths.  This inconsistency stems from using the long-wavelength (near-zero photon momentum) expansion of the electromagnetic response function, whereas the optical gyrotropy is fundamentally a finite-momentum effect (i.s. spatial dispersion of matter). Here, we resolve this long-standing problem by accurately computing the current-current optical correlation function, including all the momentum-exchange processes with the crystal lattice.
@@
where $m$ is the index of the optical band, $\varepsilon_{0}$ is the permittivity in free space, $\tilde{\bm{e}}_{m,\omega,\bm{q}}=[e_{m,\omega,\bm{q+G}_{1}},e_{m,\omega,\bm{q+G}_{2}},...]$ is the eigenvector with $e_{m,\omega,\bm{q+G}}$ as their plane-wave components, $\bm{G},\bm{G}^{\prime}$ are the reciprocal lattice vectors, and the crystal momentum $\bm{q}$ is restricted to the first Brillouin zone. The transverse SILC tensor $\bm{\sigma}^{T}_{G,G^{\prime}}(\bm{q},\omega)$ encodes the current-current correlations including sub-nanoscale momentum transfer processes within the Te crystal lattice (Supplementary).
@@
where $\bm{g}_{33}$ is the component of the optical gyrotropy tensor along the screw axis. Our theoretical plot of the optical gyrotropy in Figure 2(b) show excellent agreement with the experimental values for optical gyrotropy in Figure 2(a), providing strong validation for our theory. For the experimental values, we use the relation from [14]: $\bm{g}_{33}(\omega)\cdot\bm{q}_{z}=\Theta\cdot 2cn_{o}/\omega$, where $\Theta$ is the optical rotatory power (optical rotation per unit length) measured from experiments, and $n_{o}$ is the ordinary refractive index of Te at $\omega$ [15, 16].
@@
$\vec{\bm{e}}_{m,\bm{q}}(\bm{r},\omega)=\sum_{\bm{G}}{e}_{m,\omega,\bm{q},\bm{G}}\bm{n}^{\perp}_{\bm{q+G}}e^{i((\bm{q+G})\cdot\bm{r}-\omega t)},$ (3)
@@
[…]
@@
[…]
Experimental and simulation methods:– We pass a broadband infrared light through a linear polarizer, incident on a right-handed Te crystal. Te rotates each wavelength of light in the broadband signal by a different amount. Another linear polarizer then filters the rotated light from the crystal, and a spectrometer detects it, providing us with a correlation between the optical rotatory power and the wavelength of light.
We generate the electronic and spin-resolved optical bandstructures using our open-source software: Purdue-PicoMax [8]. We utilize the empirical pseudopotentials from [23] to generate the semi-classical electronic bandstructure, which reflects a bandgap of $0.416$eV at the H symmetry point (experimental value: 0.325eV (Supplementary)). For the optical bandstructure calculations, we use Euler method integration over the full Brillouin zone with a grid of $11\times 11\times 1$ 11 points in $k-$space, with an energy cut over the plane waves at 300eV. For the optical responses at $q=(2\pi/a)(0,0,0.1)$, we use a $k-$space grid of $15\times 15\times 15$ points, and energy cut over the plane waves at 300eV. We use gaussian averaging to remove numerical errors.

### v3

@@ Abstract
The absence of inversion symmetry in chiral tellurium (Te) creates unique spin textures within its electron waves. However, understanding textured optical waves within Te remains a challenge due to the semi-classical limitations of the long-wavelength approximation. To unveil textured optical eigenwaves in Te, we develop a spin-resolved deep-microscopic optical bandstructure, analogous to its electronic counterpart. We demonstrate that the degeneracies in this optical bandstructure is lifted by the twisted lattice of Te, which induces optical gyrotropy. Our theory shows excellent agreement with experimental optical gyrotropy measurements. At the lattice level, we reveal that the chirality of Te manifests as deep-microscopic optical spin texture within the optical wave. Our framework uncovers the finite-momentum origin of optical activity and provides a microscopic basis for light-matter interactions in chiral crystalline materials.
@@
Figure 2: Super-dispersive optical gyrotropy and chirality density for undoped Te. (a) Experimental plot of the projection of the longitudinal gyration tensor component $\bm{g}_{33}$ on the photon momentum, vs. (b) Theoretical prediction using Purdue-PicoMax at $\bm{q}=(0,0,0.1)2\pi/a$ . The Feynman diagram in the inset reflects the transverse SILC of the induced transverse current including the momentum-exchange processes. Here, $c$ and $v$ represent the conduction and valence bands respectively. (c,d) The spin-resolved chirality density projection and total spin on the isofrequency surface of RCP (++) at $\omega=60$THz (c) and LCP (-) (d). The concentration near the pole of the sphere and the difference between RCP and LCP illustrates the suitable axis (optic axis) for observing high optical activity.
@@
Figure 4: Deep-microscopic optical spin texture reveals orientation of polarization of the optical waves in Figure 3. Vectorial distribution of the optical spin texture of the optical waves in the first transverse optical band. The colormap shows the projection of the spin density along the screw axis. The plots are shown for three different planes of the Te crystal: (a) $x=0$, (b) $z=0$, and (c) $y=0$.
Our work reveals that the lack of inversion symmetry in Te lifts the degeneracy in the optical bandstructure, giving rise to optical gyrotropy. The deep-microscopic optical bandstructure captures the super-dispersive nature of this optical gyrotropy, and shows excellent agreement with experimental observations. We have implemented the necessary tools for calculating this optical gyrotropy within our open-source software: Purdue-PicoMax [8]. Previously, there have been attempts to compute the optical gyrotropy through first-principles calculations [9, 10, 11, 12, 13]. However, they do not quantitatively capture the experimentally measured dispersion of optical gyrotropy across all wavelengths. This inconsistency stems from using the long-wavelength (near-zero photon momentum) expansion of the electromagnetic response function, whereas the optical gyrotropy is fundamentally a finite-momentum effect (i.s. spatial dispersion of matter). Here, we resolve this long-standing problem by accurately computing the current-current optical correlation function, including all the momentum-exchange processes with the crystal lattice (refer Supplementary for benchmarks).
@@
where $m$ is the index of the optical band, $\varepsilon_{0}$ is the permittivity in free space, $\tilde{\bm{e}}_{m,\omega,\bm{q}}=[e_{m,\omega,\bm{q+G}_{1}},e_{m,\omega,\bm{q+G}_{2}},$...] is the eigenvector with $e_{m,\omega,\bm{q+G}}$ as their plane-wave components, $\bm{G},\bm{G}^{\prime}$ are the reciprocal lattice vectors, and the crystal momentum $\bm{q}$ is restricted to the first Brillouin zone. The transverse SILC tensor $\bm{\sigma}^{T}_{G,G^{\prime}}(\bm{q},\omega)$ encodes the current-current correlations including sub-nanoscale momentum transfer processes within the Te crystal lattice (Supplementary). We note that our theory of deep-microscopic optical bandstructure is universally applicable to all crystalline systems.
@@
where $\bm{g}_{33}$ is the component of the optical gyrotropy tensor along the screw axis. Our theoretical plot of the optical gyrotropy in Figure 2(b) show excellent agreement with the experimental values for optical gyrotropy in Figure 2(a), providing strong validation for our theory. For the experimental values, we use the relation from [14]: $\bm{g}_{33}(\omega)\cdot\bm{q}_{z}=\Theta\cdot 2cn_{o}/\omega$, where $\Theta$ is the optical rotatory power (optical rotation per unit length) measured from experiments, and $n_{o}$ is the ordinary refractive index of Te at $\omega$ [15, 16]. We note that although we extract the gyration tensor from the small $q$ limit of the full, nonlocal current-current correlation function; however, the spin-polarized optical indicators, are calculated for all momenta across the first Brillouin zone as shown in Figure 1(f).
@@
$\vec{\bm{e}}_{m,\bm{q}}(\bm{r},\omega,t)=\sum_{\bm{G}}{e}_{m,\omega,\bm{q},\bm{G}}\bm{n}^{\perp}_{\bm{q+G}}e^{i((\bm{q+G})\cdot\bm{r}-\omega t)},$ (3)
@@
where $\vec{a}_{m,\bm{q}}(\bm{r},\omega)$ is the deep-microscopic vector potential corresponding to the optical wave $\vec{e}_{m,\bm{q}}(\bm{r},\omega)$. We define the deep-microscopic optical spin texture as the spatial distribution of this optical spin density. In Figure 4, we present the deep-microscopic optical spin texture along the (a) $x=0$, (b) $z=0$, and (c) $y=0$ planes inside the Te crystal lattice for the first optical band. From Movie 2 in (Supplementary), we find that the optical spin vectors of the hidden optical waves rotate along the Te screw axis ($c=5.929\mathring${A}), within the crystal lattice.
@@
The deep-microscopic optical spin texture is a fundamental, lattice-scale quantity whose spatial average is directly related to the measurable DoCP of the hidden optical wave. The differential response of Te to varied DoCP manifests as optical gyrotropy, which serves as the key macroscopic observable of the microscopic spin texture. Although experimentally resolving this sub-lattice-level texture is challenging with the current state-of-the-art methods, we propose its direct observation using Spin-resolved Near-field Scanning Optical Microscopy (s-NSOM). This technique uses an AFM tip to scan the Te surface, interacting with and scattering the evanescent hidden optical waves. The scattered light can then be analyzed for its constituent Stokes parameters to reconstruct the full optical spin texture.
@@
The deep-microscopic optical modes with their rich spin textures offer potential avenues for dense information encoding and transfer, particularly for designing on-chip photonic waveguides. The optical gyrotropy we model is a finite-momentum effect fundamentally linked to the nonzero Berry curvature of the Bloch states [23]. While this topological origin exists in reciprocal space, our theory reveals the real-space manifestation of these optical waves and their spin at the sub-lattice level. Future work is directed towards capturing the real-space signatures of such topological features, for instance, the optical spin texture near one of the Weyl points of Te.
@@
Structural properties of Te– Te crystal is composed of helical chains of covalently bonded Te atoms along $\hat{n}_{OA}$, as shown in Figure 1(a,b). Each unit cell (highlighted in yellow) contains three Te atoms, with lattice parameters $a=4.457\mathring${A} and $c=5.929\mathring${A}. The helical chains are linked through weak van der Waals interactions. They have two different orientations or handedness, and these enantiomers are classified as the $P3_{1}21$ (right-handed) and $P3_{2}21$ (left-handed) space groups (Figure 1(c)). in this Letter, we primarily focus on the $P3_{1}21$ space group. We note that all chirality-dependent observables are perfectly inverted for $P3_{2}21$ with respect to $P3_{1}21$. For example, optical gyrotropy and the optical spin texture for the $P3_{2}21$ structure is identical to the $P3_{1}21$ structure but with the opposite sign.